OpenStudy (anonymous):

A current of 2 A is passed through a coil of resistance 75 ohm for 2 minutes,(a)how much heat energy is produced/ (b)how much charge is passed through the resistance? @butterflydreamer @perl

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the requested heat energy E is given by the subsequent formula: \[\Large E = R{I^2}\Delta t = 75 \times {2^2} \times 120 = ...Joules\]

OpenStudy (michele_laino):

since 2 mins= 120 seconds

OpenStudy (anonymous):

75 x 4 = 300?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

36k?

OpenStudy (michele_laino):

yes! 36,000 Joules

OpenStudy (anonymous):

And the next part?

OpenStudy (michele_laino):

it is simple, since the requestes amount Q of charge is equal to: \[\Large Q = I\Delta t = 2 \times 120 = ...Coulombs\]

OpenStudy (michele_laino):

requested*

OpenStudy (anonymous):

240 C.

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Help me more?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Calculate the current through a 60 W lamp rated for 250 Volts.If the voltage line falls to 200 Volts,how is the power consumed by the lamp affected?

OpenStudy (anonymous):

Current = 0.24 A?

OpenStudy (michele_laino):

here, we have to apply the subsequent formula: \[\Large P = VI\] where P is the electrical power. So the requested current is: \[\Large I = \frac{P}{V} = \frac{{60}}{{250}} = ...amperes\]

OpenStudy (anonymous):

0.24 Ampheres...

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Next part?

OpenStudy (michele_laino):

when the voltage is 250 volts, the resistance of our lamp is: \[\Large R = \frac{{{V^2}}}{P} = \frac{{{{250}^2}}}{{60}} = ...ohms\]

OpenStudy (michele_laino):

since: \[\Large P = \frac{{{V^2}}}{R}\] is another formula for electrical power

OpenStudy (anonymous):

ohms???? They what the answer in Watts.....

OpenStudy (michele_laino):

R is the electrical resistance of our lamp

OpenStudy (anonymous):

Wait, "how is the power consumed by the lamp affected?"

OpenStudy (anonymous):

Power is in Watts...

OpenStudy (michele_laino):

yes! power is measured with watts, nevertheless our lamp has an electrical resistance

OpenStudy (anonymous):

But we want the power not the resistance...

OpenStudy (michele_laino):

I know it, please follow my steps

OpenStudy (anonymous):

Ok so 1041.6 ohm?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now if we apply a voltage of 200 volts, the new current is: \[\Large {I_1} = \frac{{{V_1}}}{R} = \frac{{200}}{{1041.6}} = ...amperes\]

OpenStudy (anonymous):

0.192 A

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

and the new absorbed power is: \[\Large {P_1} = {I_1}{V_1} = 0.192 \times 200 = ...watts\]

OpenStudy (anonymous):

38.4 W

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

THnkas. Help me more??

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

An electric bulb is rated at 220V,100V Calculare:Resistance ans safe current?

OpenStudy (michele_laino):

I think: "An electric bulb is rated at 220V,100W" am I right?

OpenStudy (anonymous):

Yes, My typos t_t

OpenStudy (michele_laino):

ok! the requested resistance is: \[\Large R = \frac{{{V^2}}}{P} = \frac{{{{200}^2}}}{{100}} = ...ohms\]

OpenStudy (michele_laino):

oops.. \[\Large R = \frac{{{V^2}}}{P} = \frac{{{{220}^2}}}{{100}} = ...ohms\]

OpenStudy (anonymous):

484 ohm

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

the safe current values are given by the subsequent condition: \[\Large I \leqslant \frac{P}{V} = \frac{{100}}{{220}} = ...amperes\]

OpenStudy (anonymous):

That symbol after I ???

OpenStudy (michele_laino):

it means "less or equal to"

OpenStudy (anonymous):

0.45

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Help me more and thanks?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

A bulb rated 12V,24W rans for 20 mins Calculate: energy consumed

OpenStudy (michele_laino):

the requested energy E, is given by the subsequent computation: \[\Large E = P\Delta t = 24 \times 20 \times 60 = ...Joules\] since: \[\Large \Delta t = 20 \times 60{\text{ seconds}}\]

OpenStudy (anonymous):

28,800J

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thanks and I still have about 5 more questions can u help?

OpenStudy (michele_laino):

ok! Please post in Physics section

OpenStudy (anonymous):

Ok!